in an open pipe of length 60mm, for a complete harmonics at a velocity 330m/s, find the frequency of the waves and its wavelength if it os reduced to 40% of its original length find the wave lenght


1
Expert's answer
2021-03-08T07:31:53-0500
"f=\frac{v}{2L}=\frac{330}{2(0.06)}=2750\ Hz\\\lambda=2L=12\ cm\\\lambda'=2(0.4)L=4.8\ cm"